# Automatic Reviewers Fail to Detect Faulty Reasoning in Research Papers: A New Counterfactual Evaluation Framework

### v1

@@ Acknowledgements
This work has been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). This work has been co-founded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science, and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. This research work has been co-funded by the European Union (ERC, InterText, 101054961). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ A.3.1 Prompts
The generation of counterfactuals involves multiple steps and prompts including self-refinement. Due to space restrictions, we report on the exact prompts used for counterfactual generation in the supplementary code.

### v2

@@ 7 Limitations and Future Work
In this Section, we summarize potential limitations of the proposed framework and point to important future work.
Assumptions
@@
Interpretation
Our experiments show that state-of-the-art ARGs fail to identify flawed research logic in research papers. We further find that ARGS are sensitive to surface-level features, suggesting that the underlying LLMs rely on superficial heuristics instead of thorough reasoning. Investigating the causes of this behavior is an important direction for future work. In particular, distinguishing failures to identify soundness-relevant information in the paper from failures to recognize logical fallacies within the research logic itself are crucial for developing more robust ARGs in the future.
Reproducibility
For dataset creation and evaluation we primarily employ closed commercial models, as their outputs demonstrated higher quality during manual tuning. To enhance replicability, we release all model outputs in the supplementary materials. For reproducibility, detailed model versions are reported in Appendix B.2. However, the use of closed models inevitably limits future reproduction on new data, as such models may undergo unrecorded changes. In future work, the extension to open models that perform better or on-par with commercial LLMs for these steps is a promising avenue to enhance reproducibility.
@@ Acknowledgements
This work has been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). This work has been co-founded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science, and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. This research work has been co-funded by the European Union (ERC, InterText, 101054961). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. We express our sincere gratitude to the reviewers and action editor at TACL for their valuable and constructive feedback.
@@ A.3.1 Prompts
The generation of counterfactuals involves multiple steps and prompts including self-refinement. Due to space restrictions, we report on the exact prompts used for counterfactual generation in the supplementary code. Figure 6 summarizes the general prompt structure which is common to all steps in the pipeline.
[system persona prompt]
[brief summary of the task]
### DEFINITIONS [definition of terms and concepts]
### INPUTS [structured list of inputs]
### SPECIFIC INSTRUCTIONS [step-by-step instructions]
### OUTPUT FORMAT [description of output JSON format]
### IMPORTANT POINTS [reiterating hard constraints and requirements]
Figure 6: General prompt design used during dataset creation.